OpenStudy (31356):

(3x)^3 Show your work to apply the laws of exponents! FREE MEDAL REWARDED!!

jimthompson5910 (jim_thompson5910):

hint: use the idea that (a*b)^x = a^x * b^x

OpenStudy (31356):

I know, just show the work,please? I am not that good at Math, or else I wouldn't be here

jimthompson5910 (jim_thompson5910):

Similar Example (5x)^2 = 5^2*x^2 (5x)^2 = 25x^2

jimthompson5910 (jim_thompson5910):

Another Similar Example (7w)^4 = 7^4*w^4 (7w)^4 = 2401w^4

OpenStudy (31356):

Thanks!

jimthompson5910 (jim_thompson5910):

you're welcome